[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Nov 07, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Nov 8 03:03:47 UTC 2017


HSU Shift Log for Tue Nov 07 2017 22:03:41 GMT-0500 (EST)

Shift: R

Forecaster: D Brown

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: Few missing G-16 images around 21-22z, but none since.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171108/ccba939b/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list